OpenStudy (anonymous):

Solve: (3sqrt2-9sqrt3)(5sqrt2+15sqrt3) Please help

OpenStudy (radar):

\[(3\sqrt{2}-9\sqrt{3})(5\sqrt{2}+15\sqrt{3})\]

OpenStudy (anonymous):

yes that is the problem...do you know how to do it?

OpenStudy (radar):

Multiply using rules that you need to know (15)(2)-45sqrt6+45sqr6-(135)(3) simplify: 30-405=-375

OpenStudy (radar):

Do you know the rules for multiplying radicals?

OpenStudy (anonymous):

Somewhat

OpenStudy (anonymous):

This class is killing me so bad....confusing

OpenStudy (radar):

Take your time and practice

OpenStudy (anonymous):

Thank you so much

OpenStudy (radar):

\[\sqrt{x}\sqrt{y}=\sqrt{xy}\]

OpenStudy (anonymous):

That makes sense. but why did you pull the 2 out of the matching radicals

OpenStudy (radar):

\[\sqrt{x}\sqrt{x}=x\]

OpenStudy (anonymous):

Awwww...that's where i was making my mistake then...thank you so much for the help!

OpenStudy (radar):

\[\sqrt{2}\sqrt{2}=2\]